OpenStudy (anonymous):

reduce 16/75 to the lowest terms

Parth (parthkohli):

I don't see it reducible.

Parth (parthkohli):

Irreducible in other words.

OpenStudy (anonymous):

i cant see it either

Parth (parthkohli):

Lol who gave such a question to you

OpenStudy (accessdenied):

Yeah, the numerator factors only into 2's, and the denominator is odd; so nothing can be reduced. :P

OpenStudy (anonymous):

my brother just to see if it was reducible

OpenStudy (accessdenied):

I personally find the best way to tell whether it is reducible is to break it into the prime factors. If there's any prime factors of the numerator also in the denominator, then they can be taken out. 16 = 4*4 = 2*2*2*2 75 = 3*25 = 3*5*5 ^ Nothing is shared, so it's not reducible. If we had like 25/75... 25 = 5*5 75 = 3*5*5 We can just take out the two 5's from numerator/denominator: 1/3.

OpenStudy (anonymous):

another question what is the answer to 24 divided into 18478

OpenStudy (accessdenied):

Hmm... 18478/24 They're both even, so we can take out a 2/2 9329/12 From here, we cannot factor out either a 2 or a 3, the two unique factors of 12, so it looks like this is the simplest form of the fraction.

OpenStudy (accessdenied):

9239 ***, not 9329

OpenStudy (anonymous):

thanks

OpenStudy (accessdenied):

You're welcome. :) 9239/12 should be correct, just to make sure.